# On the hypothesis of a bi-isotropic plasma permeating the interstellar space

### v1

@@ IV.1 Dispersion measure and rotation measure in chiral interstellar medium
$\left(v_{g}\right)_{L,R}^{-1}\approx 1\pm\tilde{\xi_{c}}+\frac{\omega_{p}^{2}}{2\omega^{2}},$ (36)
@@ IV.2 Constraints using LOFAR dataset
Observational dispersion measure values for several pulsars are available in the LOFAR census dataset [62], from which, for our estimates, we have selected five, namely, B1919+21, B1944+17, B1929+10, B2016+28, and B2020+28. Uncertainties associated with each measure, here denoted by $\epsilon_{\text{DM}}$, are also available in the catalog and will be assigned to the chiral parameter in our present proposal. Although it may consist of a non-rigorous approach, it can be employed to provide a first constraining estimate, justified by the assumption of a small chiral parameter in the bi-isotropic interstellar plasma. Therefore, for each measurement, we can constrain the magnitude of the chiral correction (38) by the corresponding DM uncertainty, that is,
@@ V Final remarks
In this work, we have investigated the propagation of electromagnetic waves in a cold and magnetized bi-isotropic chiral plasma by discussing optical properties and establishing upper bounds on the chiral parameter magnitude. The bi-isotropic constitutive relations allowed the derivation of four modified refractive indices given in Eqs. (18) and (19), associated with RCP and LCP propagating modes, respectively. The chiral term promotes significant modifications to the electromagnetic propagation inside this plasma medium. For instance, in Sec. II, negative refraction behavior appears in the range $\omega_{c}<\omega<\omega_{-}$ for the index $n_{R_{+}}$ (see Fig. 1), with absorption ($\mathrm{Im}[n_{R_{+}}]\neq 0$) occurring for $\omega_{c}<\omega<\omega_{R+}$ and free (metamaterial-like) propagation in the range $\omega_{R+}<\omega<\omega_{-}$. The existence of negative refraction in a chiral plasma has already been investigated and represents a typical feature of bi-isotropic plasmas [70, 69]. Moreover, in the low-frequency regime, modified propagating RCP helicons arise exhibiting a linear dependence in the chiral parameter $\tilde{\xi_{c}}$, as discussed in Subsection. II.1.
@@
By supposing a finite chiral parameter, we study optical effects expected to occur in the cold plasma system, such as optical rotation and dichroism. In this sense, we have discussed its consequences through both the RP and dichroism coefficients, which were determined starting from the refractive indices $n_{R_{+}}$, $n_{L_{+}}$ and $n_{L_{-}}$ (see Sec. III). We highlight that the RP $\delta_{LR}^{++}$ defined by Eq. (30), depicted in Fig. 5, presents a double sign reversal at the frequencies $\omega=\omega^{\prime}$ and $\omega=\omega^{\prime\prime}$, located in the range $0<\omega<\omega_{c}$. For $\omega>\omega_{c}$, the RP is negative and decreases with the frequency, as shown in Fig. 6. Likewise, the double reversal also appears in interval $0<\omega<\omega_{+}$ for the RP $\delta_{LR}^{-+}$ given in (32), as displayed in Fig. 7. Although cold plasmas do not exhibit such a double RP reversal, a single reversion already was reported in both rotating plasmas [80] and cold chiral plasma ruled by the axion electrodynamics [66, 67]. Notwithstanding being a non-common optical signature, the double RP reversal has also been encountered in bi-isotropic media with anomalous Hall current [82]. This peculiar birefringent property may represent a route for an optical characterization of the bi-isotropic chiral plasmas in realistic experimental setups.
@@
Similarly, Subsec. III.2 examines the effects of the chiral term on the dichroism phenomenon through coefficients $\delta_{dRL}^{++}$ and $\delta_{dLR}{-+}$, which are non-null only in the intervals $0<\omega<\omega_{R+}$ and $\omega_{c}<\omega<\omega_{R-}$, such as illustrated in Fig. 8. Compared to the standard case, the chiral parameter of bi-isotropic plasma narrows the zones where dichroism occurs, meaning that attenuation-free propagation of larger frequency bands is now allowed.
@@
In order to conciliate the hypothesis of a bi-isotropic plasma permeating the interstellar space with available observational data, we have addressed an ISM plasma model ruled by the Maxwell electrodynamics endowed with the constitutive relations (12), yielding the modified refractive indices discussed in Sec. II. Thus, in Sec. IV, we have utilized astrophysical data of dispersion and rotation measures obtained from radio signals traveling in a low-density ionized medium to analyze the implications of the chiral parameter in the electromagnetic propagation. Such effects have allowed us to find upper bounds constraining the chiral variable. From the time delay (37), a chiral dispersion measure contribution, $\text{DM}_{chiral}$, was obtained and limited to uncertainty measures ($\epsilon_{\text{DM}}$) according to (43). Analogously, a chiral rotation measure, $\text{RM}_{chiral}$, was obtained from the Faraday rotation (40), and limited to uncertainty $\epsilon_{\text{RM}}$ in (44). Similarly, we impose astrophysical limits on the chiral parameter by adopting data of dispersion and rotation measures attained for five pulsars: B1919+21, B1944+17, B1929+10, B2016+28, and B2020+28. Then, by using dispersion measure (DM) data, the $\tilde{\xi}$ magnitude was restricted to the orders between $10^{-16}$$10^{-18}$, as presented in the fourth column of Table 1. Furthermore, utilizing rotation measure (RM) information, the chiral factor was restrained as $\tilde{\xi}\lesssim 10^{-22}$$10^{-21}$ (see the last column in Table 1).
@@ Acknowledgements.
The authors express their gratitude to FAPEMA, CNPq, and CAPES (Brazilian research agencies) for their invaluable financial support. M.M.F. is supported by FAPEMA APP-12151/22, CNPq/Produtividade 317048/2023-6 and CNPq/Universal/422527/2021-1. P.D.S.S. is grateful to FAPEMA APP-12151/22. R. C. acknowledges the support from the grants CNPq/312155/2023-9, FAPEMA/UNIVERSAL 00812/19, and FAPEMA APP-12299/22. Furthermore, we are indebted to CAPES/Finance Code 001 and FAPEMA/POS-GRAD-04755/24.

### v2

@@ IV.1 Dispersion measure and rotation measure in chiral interstellar medium
$\left(v_{g}\right)_{L,R}^{-1}\approx\frac{1}{c}\pm\frac{\tilde{\xi_{c}}}{c}+\frac{\omega_{p}^{2}}{2\omega^{2}},$ (49)
@@ IV.2 Constraints using LOFAR dataset
Observational dispersion measure values for several pulsars are available in the LOFAR census dataset [62], from which, for our estimates, we have selected five, namely, B1919+21, B1944+17, B1929+10, B2016+28, and B2020+28. Uncertainties associated with each measure, here denoted by $\epsilon_{\text{DM}}$, are also available in the catalog and will be assigned to the chiral parameter in our present proposal. Although it may consist of a nonrigorous approach, it can be employed to provide a first constraining estimate, justified by the assumption of a small chiral parameter in the bi-isotropic interstellar plasma. Therefore, for each measurement, we can constrain the magnitude of the chiral correction (51) by the corresponding DM uncertainty, that is,
@@ V Final remarks
In this work, we have investigated the propagation of electromagnetic waves in a cold and magnetized bi-isotropic-like chiral plasma by discussing optical properties and establishing upper bounds on the chiral parameter magnitude. The bi-isotropic constitutive relations allowed the derivation of four modified refractive indices given in Eqs. (31) and (32), associated with RCP and LCP propagating modes, respectively. The chiral term promotes significant modifications to the electromagnetic propagation inside this plasma medium. For instance, in Sec. II, negative refraction behavior appears in the range $\omega_{c}<\omega<\omega_{-}$ for the index $n_{R_{+}}$ (see Fig. 1), with absorption ($\mathrm{Im}[n_{R_{+}}]\neq 0$) occurring for $\omega_{c}<\omega<\omega_{R+}$ and free (metamaterial-like) propagation in the range $\omega_{R+}<\omega<\omega_{-}$. The existence of negative refraction in a chiral plasma has already been investigated and represents a typical feature of bi-isotropic plasmas [95, 94]. Moreover, in the low-frequency regime, modified propagating RCP helicons arise exhibiting a linear dependence in the chiral parameter $\tilde{\xi_{c}}$, as discussed in Sec. II.1.
@@
By supposing a finite chiral parameter, we study optical effects expected to occur in the cold plasma system, such as optical rotation and dichroism. In this sense, we have discussed its consequences through both the RP and dichroism coefficients, which were determined starting from the refractive indices $n_{R_{+}}$, $n_{L_{+}}$ and $n_{L_{-}}$ (see Sec. III). We highlight that the RP $\delta_{LR}^{++}$ defined by Eq. (43), depicted in Fig. 5, presents a double sign reversal at the frequencies $\omega=\omega^{\prime}$ and $\omega=\omega^{\prime\prime}$, located in the range $0<\omega<\omega_{c}$. For $\omega>\omega_{c}$, the RP is negative and decreases with the frequency, as shown in Fig. 6. Likewise, the double reversal also appears in interval $0<\omega<\omega_{+}$ for the RP $\delta_{LR}^{-+}$ given in (45), as displayed in Fig. 7. Although cold plasmas do not exhibit such a double RP reversal, a single reversion already was reported in both rotating plasmas [106] and cold chiral plasma ruled by the axion electrodynamics [91, 92]. Notwithstanding being a noncommon optical signature, the double RP reversal has also been encountered in bi-isotropic media with anomalous Hall current [108]. This peculiar birefringent property may represent a route for an optical characterization of the bi-isotropic chiral plasmas in realistic experimental setups.
@@
Similarly, Sec. III.2 examines the effects of the chiral term on the dichroism phenomenon through coefficients $\delta_{dRL}^{++}$ and $\delta_{dLR}{-+}$, which are non-null only in the intervals $0<\omega<\omega_{R+}$ and $\omega_{c}<\omega<\omega_{R-}$, such as illustrated in Fig. 8. Compared to the standard case, the chiral parameter of bi-isotropic plasma narrows the zones where dichroism occurs, meaning that attenuation-free propagation of larger frequency bands is now allowed.
@@
In order to conciliate the hypothesis of a bi-isotropic-like plasma permeating the interstellar space with available observational data, we have addressed an ISM plasma model ruled by the Maxwell electrodynamics endowed with the constitutive relations (23), yielding the modified refractive indices discussed in Sec. II. Thus, in Sec. IV, we have utilized astrophysical data of dispersion and rotation measures obtained from radio signals traveling in a low-density ionized medium to analyze the implications of the chiral parameter in the electromagnetic propagation. Such effects have allowed us to find upper bounds constraining the chiral variable. From the time delay (50), a chiral dispersion measure contribution, $\text{DM}_{chiral}$, was obtained and limited to uncertainty measures ($\epsilon_{\text{DM}}$) according to (56). Analogously, a chiral rotation measure, $\text{RM}_{chiral}$, was obtained from the Faraday rotation (53), and limited to uncertainty $\epsilon_{\text{RM}}$ in (57). Similarly, we impose astrophysical limits on the chiral parameter by adopting data of dispersion and rotation measures attained for five pulsars: B1919+21, B1944+17, B1929+10, B2016+28, and B2020+28. Then, by using dispersion measure (DM) data, the $\tilde{\xi}$ magnitude was restricted to the orders between $10^{-16}$$10^{-18}$, as presented in the fourth column of Table 1. Furthermore, utilizing rotation measure (RM) information, the chiral factor was restrained as $\tilde{\xi}\lesssim 10^{-22}$$10^{-21}$ (see the last column in Table 1).
@@ Acknowledgements.
The authors express their gratitude to FAPEMA, CNPq, and CAPES (Brazilian research agencies) for their invaluable financial support. M.M.F. is supported by Grants No. FAPEMA APP-12151/22, No. CNPq/Produtividade 317048/2023-6, and No. CNPq/Universal/422527/2021-1. P.D.S.S. is grateful to FAPEMA APP-12151/22. R. C. acknowledges the support from the Grants No. CNPq/312155/2023-9, No. FAPEMA/UNIVERSAL 00812/19, and No. FAPEMA APP-12299/22. Furthermore, we are indebted to CAPES/Finance Code 001 and FAPEMA/POS-GRAD-04755/24.
DATA AVAILABILITY
The data that support the findings of this article are openly available in Refs. [62, 63, 109].